OpenStudy (anonymous):

The included angle of the two sides of constant equal length s of an isosceles triangle is Ө. a. Show that the area of the triangle is given by A= 1/2s^2sinӨ b. If Ө is increasing at the rate of 1/2 radians per minute, find the rates of change of the area when Ө is when Ө=π/6 and Ө=π/3. c. Explain why the rate of change of the area of the triangle is not constant even though dӨ/dt is constant. PLEASE HELP GUYS, THIS IS HALF OF MY WHOLE GRADE.

OpenStudy (anonymous):

For a) you need to draw the triangle: Label the height H and the base B and then you know that looking at the right triangle formed between one of the sides of S and the height line (draw the height line) is \[ \sin{\theta}=\frac{opp.}{hyp.}=\frac{H}{S}\] and that \[\cos{\theta} = \frac{adj.}{hyp.} = \frac{\frac{1}{2}B}{S}\] Since the area is \[\frac{1}{2}BH\] you need to solve for B and H then sub. them into the area formula

OpenStudy (anonymous):

Oh alright, I get it thank you that's one step down :D

OpenStudy (anonymous):

Wait, instead of theta you need to use (360-theta)/2

OpenStudy (anonymous):



OpenStudy (anonymous):

I see but how do you solve for B and H and then substitute them into the area formula?

OpenStudy (anonymous):

Grrr, I'm still confused. I am so not good at these types of problems -.-

OpenStudy (anonymous):

Like the only way I understand if by viewing how the WHOLE problem is done, it's how I've been teached.

OpenStudy (anonymous):

 I am going to call (360-theta)/2 "Z" for now and the resub later. 1) from SOHCAHTOA, sin(z) = H/S, therefore H = (S)sin(z) 2) from SOHCAHTOA cos(z) = 0.5B/S, therefore B = (2)(S)cos(z) 3) Area = 0.5(B)(H) = 0.5 (2)(S)(cos(z)) (S)sin(z) The only issue now is that you have to use the sin(a+b) and cos(a+b) to finish where z= (180+theta/2) sin(a+b) = sin(a)cos(b)-cos(a)sin(b) cos(a+b) = cos(a)cos(b)-sin(a)sin(b) You will also need to remember that 2sin(a)cos(a) = sin(2a)

OpenStudy (anonymous):

and that cos(-x) = cos(x) and that sin(-x) = -sin(x)

OpenStudy (anonymous):

Sorry z = 180-(theta/2) mis typed.

OpenStudy (anonymous):

For example \[sin{\left(\pi -\frac{\theta}{2}\right)}=\sin{(\pi)}\cos{\left(-\frac{\theta}{2}\right)}-\cos{(\pi)}\sin{\left(-\frac{\theta}{2}\right)}\] since sin(pi) = 0 \[sin{\left(\pi -\frac{\theta}{2}\right)}=0-\cos{(\pi)}\sin{\left(-\frac{\theta}{2}\right)}\] since cos(pi) = -1 \[sin{\left(\pi -\frac{\theta}{2}\right)}=+\sin{\left(-\frac{\theta}{2}\right)}\] since sin(-x)=-sin(x) \[sin{\left(\pi -\frac{\theta}{2}\right)}=-\sin{\left(\frac{\theta}{2}\right)}\]